# Mapping the Cosmic Gamma-ray Horizon: The 1CGH Catalogue of Fermi-LAT detections above 10 GeV

### v2

@@ Abstract
gamma-rays: galaxies, catalogues – radiation mechanisms: non-thermal
@@ 1 Introduction
The attenuation of $\gamma$-rays by the EBL defines a statistical boundary for the farthest distance at which VHE photons of a given energy are likely to be detected, known as the Cosmic Gamma-ray Horizon (CGH). The concept of the CGH was initially explored by Nikishov (1961); Gould & Schréder (1967), while Fazio & Stecker (1970) -with the Fazio Stecker relation- advanced the discussion of $\gamma$-ray attenuation in context of a cosmological origin for the isotropic $\gamma$-ray background.
@@ 2.1 Selection of Gamma-ray Candidate Sources
To create a sample of sky positions and define our $\gamma$-ray targets, we combined several major blazar and AGN catalogues. We began by merging the ‘5th edition of the Roma-BZCAT Multi-Frequency Blazar Catalog’ (5BZcat, Massaro et al., 2015a) with the ‘3rd edition of the High Synchrotron Peak sample’ (3HSP, Chang et al., 2019). These catalogues were chosen for their extensive blazar coverage and reliability.
Both 5BZcat and 3HSP are advanced versions of multifrequency catalogues containing thousands of blazars and blazar candidates. These catalogues have been iteratively refined over time (e.g., refer to BZcat, 1WHSP, and 2WHSP catalogues for earlier versions Massaro et al., 2009; Arsioli et al., 2015; Chang et al., 2017), and represent significant repositories for the very high-energy community.
@@
We further include all sources classified as AGNs, blazars, and blazar candidates in the ‘Fourth Catalog of Active Galactic Nuclei Detected by the Fermi Large Area Telescope: Data Release 3’ (4LAC-DR3, Ajello et al., 2022), since many of those candidates are not covered by the 5BZcat and 3HSP catalogues.
To remove duplicate entries, we use TopCat (Taylor, 2005) for internal matching; For the 5BZcat and 3HSP samples, the R.A. and Dec. positions, corresponding to the optical and infrared counterparts respectively, were used. For TeVcat and 4LAC-DR3 sources, we used the associated counterpart position whenever available (e.g., radio, infrared, or optical), as these typically offer better astrometric precision compared to the $\gamma$-ray position. With this sample of unique sources, we searched for $\gamma$-ray associations in the 4FGL-DR4 catalogue (Ballet et al., 2023) and the 3FHL catalogue (Ajello et al., 2017), considering the 95% containment region reported for each source.
@@ 2.2 Broadband Likelihood Analysis with Fermi-LAT Science Tools
In our analysis, we adopt a low-energy threshold of 10 GeV –similar to the 3FHL catalogue– due to the improved angular resolution and lower background contamination compared to lower thresholds (Ajello et al., 2017). For the upper energy limit, we set a conservative threshold of 800 GeV, whereas 3FHL extended the analysis up to 2 TeV. This decision follows from the discussion in the 4FGL paper (Section 3.2 Abdollahi et al., 2020), which highlights that a broadband analysis reaching 1 TeV can introduce uncertainties in the energy flux estimate, particularly for hard-spectrum sources. Therefore, our choice for the high-energy threshold is meant to mitigate uncertainties regarding the spectral fitting of all sources in the ROI region.
@@ 2.3 Detection Threshold for Multifrequency-Selected Targets
The use of multifrequency seed positions to detect $\gamma$-ray sources has been successfully applied in numerous works (Arsioli & Chang, 2017; Arsioli et al., 2018; Arsioli & Polenta, 2018a; Arsioli et al., 2020) and has proven effective in uncovering extreme blazars close to the Fermi-LAT detection threshold. Using this approach, the analysis for each sky position is independent, investigating a single $\gamma$-ray candidate at a time. To handle all 8200 candidates efficiently, we relied on parallel High Performance Computing (HPC) resources provided by INCD-Portugal.
@@
In addition, this relatively low selection threshold aims not only to highlight faint sources but also to capture those that undergo short-lived flare episodes, which might otherwise be missed over a long integration period due to signal dilution (Arsioli & Polenta, 2018b). As the true signal (i.e. the photon counts from a transient flare) remains constant while the background noise accumulates, a lower TS threshold allows us to include those valuable cases. In short time windows, the significance of these flaring events would be higher, enabling the study of the effects of the EBL absorption. Although we did not perform light-curve analyses in this work, the 1CGH catalogue can serve as a basis for future time-resolved studies.
To assess the robustness of our detection threshold, we estimated the spurious detection rate in our analysis setup. In a chi-squared ($\chi^{2}$) distribution with two degrees of freedom (d.o.f), the cumulative probability $P(TS>TS_{threshold})$ represents the likelihood of obtaining a test statistic above the threshold purely by chance. Using TS=12, the spurious detection probability is calculated as $\rm p\_value=1-\chi^{2}\times cdf(TS=12,\ d.o.f=2)\approx 0.0025$, where $cdf()$ is the cumulative distribution function. Given that 8200 seed positions were analysed, the expected number of spurious detections in the pre-selection phase is approximately $\rm 8200\times 0.0025=20.5$. This value represents an upper limit on contamination.
@@ 2.4 Additional Selection Criteria: Refining the 1CGH Sample
In the Fermi-LAT PASS8 release (Atwood et al., 2013), each event is assigned a Point Spread Function (PSF) type, which indicates the quality of the reconstructed direction. Events are categorized into four quartiles: PSF0 (lowest quality, evtype=4), PSF1 (evtype=8), PSF2 (evtype=16), and PSF3 (highest quality, evtype=32). The PSF 68% containment radius varies by energy and event type. For instance, at 30 GeV, the containment radius for PSF1 events is approximately 0.12, as described in the PASS8 documentation [^4] Fermi-LAT PASS8 PSF: https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm.
@@ 3 The First Cosmic Gamma-ray Horizon Catalogue (1CGH)
Figure 1: The distribution of detection significance ($\sigma$) for the 1CGH and 3FHL samples (respectively, blue and red-dashed lines).
In Fig. 1 we show the distribution of detection significance for the entire 1CHG and 3FHL catalogues. This representation is meant as a qualitative overview on both samples, keeping in mind that the 1CGH is focus on seed-positions of blazars and blazar-candidates, while the 3FHL includes detections of all source types. The number of detections in 1CGH shows consistent growth along the significance-bins (i.e. following a trend similar to 3FHL), and includes an extra bin covering the 3<$\sigma$<4 range.
We should note that, in the full 1CGH catalogue, 26 sources [^8] Comprising 3 unassociated sources, 4 BL Lacs, 8 Flat Spectrum Radio Quasars (FSRQs), and 11 blazars of unknown type. reached the spectral index limit of $\Gamma=6.5$ constrained by our likelihood analysis setup (see Section 2.2). Moreover, 27 1CGH sources have counterparts in 3FHL where spectral curvature has been identified [^9] To help identify such cases, we include the Curvature_3FHL column to our final catalogue, which tracks whether spectral curvature was identified in 3FHL, and might affect broadband flux estimates.. For these 53 (26+27) cases, the derived fluxes should be interpreted as upper limits. While this affects a small fraction of the sample, the vast majority of sources exhibit well-constrained spectral fits, reinforcing the robustness of the 1CGH catalogue for high-energy studies.
@@ 3.2 The Cosmic Gamma-ray Horizon Plot
Here we should draw attention to the number of 1CGH sources which lack a proper redshift characterization: Of the total 2791 1CGH sources, 1062 (38.2%) have spectroscopically measured redshift values, 855 (30.6%) have uncertain/photometric values, 210 (7.5%) have lower limit values, and 664 (23.7%) have no available redshift information. Therefore, 1729 ($\sim$72%) of the 1CGH catalogue would benefit from dedicated observation campaigns to improve the overall redshift description.
@@ 3.3 Identification of Sources with Detectable EBL Absorption
In Figure 4, we illustrate the energy versus redshift relation and use the Saldana-Lopez EBL model Saldana-Lopez et al. (2021) as a reference framework to identify sources where gamma-ray absorption might be detectable. The Saldana-Lopez model offers an empirical determination of the evolving EBL spectrum and its uncertainties up to $\sim$6; it is based on galaxy counts and multifrequency data (from UV to far-IR) covering the Hubble Space Telescope’s CANDELS survey (Grogin et al., 2011), and was designed to minimise uncertainties especially at higher redshift.
Our ultimate goal, however, is to highlight a sample in which $\tau_{(E,z)}$ values can be derived from the data, independently of any specific EBL model. The estimates of Saldana-Lopez et al. (2021) are similar to those by Finke et al. (2010) and both are representative of the most severe attenuation models among the ones considered.
We use an EBL optical depth of $\rm\tau_{(E,z)}=0.1$ as a lower limit to introduce an absorption flag (ABS-flag), indicating whether spectral absorption might be detectable at the highest energy bins. For cases where $\rm E^{bin}_{max}$ lies in the $\rm\tau_{(E,z)}>0.1$ regime, the largest energy bin observed with Fermi-LAT likely experiences a detectable absorption greater than 10% (i.e., >($\rm 1-e^{-0.1}$)). These sources can be used to investigate the opacity along redshift, and have an ABS-flag set to ‘1’.
@@
The $\rm\tau_{(E,z)}$ values  –calculated using the Saldana-Lopez EBL model at $\rm E^{bin}_{max}$ Saldana-Lopez et al. (2021)– are listed under the column ‘$\rm\tau_{Ebin}$’. Additionally, the absorption fraction ($\rm 1-e^{-\tau_{(E,z)}}$) is recorded in the ‘$\rm Abs_{Ebin}$’ column, which can be used to create subsamples based on different absorption thresholds. Notably, for sources with multiple energy bins experiencing detectable absorption, a single source can provide $\rm\tau_{(E,z)}$ estimates for multiple energy levels at a given redshift.
@@ 4 Summary and Conclusions
In this work, we presented the First Cosmic Gamma-ray Horizon (1CGH) catalogue, which lists all-sky Fermi-LAT detections above 10 GeV, compiled from 16 years of observations. The catalogue includes 2791 sources, with 62 representing new $\gamma$-ray detections, and to date, it represents an advanced sample to study the transparency of the universe to VHE photons.
We significantly improved the redshift characterization of our catalogue by incorporating new spectroscopic redshift estimates whenever possible. By meticulously reviewing nearly 70 dedicated observational campaigns, we reduced a substantial portion of the redshift knowledge-gap, categorising redshift reliability through our z-flag system.
The 1CGH catalogue highlights sources experiencing varying degrees of $\gamma$-ray attenuation, and will contribute to a better determination of the EBL’s density and its evolution over cosmic time. Using the Saldana-Lopez EBL model as a reference framework, we identified sources with $\gamma$-ray absorption above the $\rm\tau_{(E,z)}>0.1$ limit, creating a targeted sample where absorption is likely measurable.
In addition, we introduced a complementary data-exploration method by calculating the mean energy of the four highest-energy photons $\rm E^{bin}_{max}$, rather than focusing solely on the highest-energy event. This provides a more stable and statistically reliable measure of the characteristic ‘largest energy level’ observed for each source.
We have identified sources that should be prioritised for follow-up optical observation campaigns. These include: (i) sources that lack redshift information but have a clear optical or radio association, prioritised based on the highest $\rm E^{bin}_{max}$ values; and (ii) sources with lower limit or uncertain redshift estimates that could benefit from further spectroscopic exploration, prioritised by their current $\tau$ values at $\rm E^{bin}_{max}$. These targets will not only help refine the redshift description of the catalogue but also improve the accuracy of EBL measurements at higher redshift, thus reducing uncertainties in SFR density determination during the EoR, at $\sim$6-7.
@@ Appendix A Column Descriptions of the 1CGH Catalogue
The 1CGH catalogue contains detailed metadata for 2791 sources detected above 10 GeV based on 16 years of Fermi-LAT observations. Table 2 provides a description of each column, detailing its content.
@@
zflag_3HSP Redshift flag from the 3HSP catalogue.

### v3

@@ Abstract
gamma-rays: galaxies – galaxies: active – catalogues – radiation mechanisms: non-thermal
@@ 1 Introduction
The attenuation of $\gamma$-rays by the EBL defines a statistical boundary for the farthest distance at which VHE photons of a given energy are likely to be detected, known as the Cosmic Gamma-ray Horizon (CGH). The CGH concept was initially explored by Nikishov (1961); Gould & Schréder (1967), while Fazio & Stecker (1970) -with the Fazio Stecker relation- advanced the discussion of $\gamma$-ray attenuation in context of a cosmological origin for the isotropic $\gamma$-ray background.
@@ 2.1 Selection of Gamma-ray Candidate Sources
To create a sample of sky positions and define our $\gamma$-ray targets, we combined several major blazar and AGN catalogues. We began by merging the ‘5th edition of the Roma-BZCAT Multi-Frequency Blazar Catalog’ (5BZcat; Massaro et al., 2015a) with the ‘3rd edition of the High Synchrotron Peak sample’ (3HSP; Chang et al., 2019). These catalogues were chosen for their extensive blazar coverage and reliability.
Both 5BZcat and 3HSP are advanced versions of multifrequency catalogues containing thousands of blazars and blazar candidates, and represent significant repositories for the VHE community. These catalogues have been iteratively refined over time, building upon the initial BZcat (Massaro et al., 2009), and the First and Second Catalogues of High Synchrotron Peaked Blazars and Candidates, 1WHSP and 2WHSP (Arsioli et al., 2015; Chang et al., 2017).
@@
We further include all sources classified as AGNs, blazars, and blazar candidates in the ‘Fourth Catalog of Active Galactic Nuclei Detected by the Fermi Large Area Telescope: Data Release 3’ (4LAC-DR3; Ajello et al., 2022), since many of those candidates are not covered by the 5BZcat and 3HSP catalogues.
To remove duplicate entries, we use TopCat (Taylor, 2005) for internal matching; For the 5BZcat and 3HSP samples, the R.A. and Dec. positions, corresponding to the optical and infrared counterparts respectively, were used. For TeVcat and 4LAC-DR3 sources, we used the associated counterpart position whenever available (e.g., radio, infrared, or optical), as these typically offer better astrometric precision compared to the $\gamma$-ray position. With this sample of unique sources, we searched for $\gamma$-ray associations in the ‘Fermi LAT 14-Year Point Source Catalogue’ (4FGL-DR4; Ballet et al., 2023) and the 3FHL catalogue (Ajello et al., 2017), considering the 95% containment region reported for each source.
@@ 2.2 Broadband Likelihood Analysis with Fermi-LAT Science Tools
In our analysis, we adopt a low-energy threshold of 10 GeV –similar to the 3FHL catalogue– due to the improved angular resolution and lower background contamination compared to lower thresholds (Ajello et al., 2017). For the upper energy limit, we set a conservative threshold of 800 GeV, whereas 3FHL extended the analysis to 2 TeV. This decision follows from the discussion in the 4FGL paper (Section 3.2 of Abdollahi et al., 2020), which highlights that a broadband analysis reaching 1 TeV can introduce uncertainties in the energy flux estimate, particularly for hard-spectrum sources. Therefore, our choice for the high-energy threshold is meant to mitigate uncertainties regarding the spectral fitting of all sources in the ROI region.
@@ 2.3 Detection Threshold for Multifrequency-Selected Targets
The use of multifrequency seed positions to detect $\gamma$-ray sources has been successfully applied in numerous works (Arsioli & Chang, 2017; Arsioli et al., 2018; Arsioli & Polenta, 2018; Arsioli et al., 2020) and has proven effective in uncovering extreme blazars close to the Fermi-LAT detection threshold. Using this approach, the analysis for each sky position is independent, investigating a single $\gamma$-ray candidate at a time. To handle all 8200 candidates efficiently, we relied on parallel High Performance Computing (HPC) resources provided by INCD-Portugal.
@@
In addition, this relatively low selection threshold aims not only to highlight faint sources but also to capture those that undergo short-lived flare episodes, which might otherwise be missed over a long integration period due to signal dilution (Arsioli & Polenta, 2018). As the true signal (i.e. the photon counts from a transient flare) remains constant while the background noise accumulates, a lower TS threshold allows us to include those valuable cases. In short time windows, the significance of these flaring events would be higher, enabling the study of the effects of the EBL absorption. Although we did not perform light-curve analyses in this work, the 1CGH catalogue can serve as a basis for future time-resolved studies.
To assess the robustness of our detection threshold, we estimated the spurious detection rate in our analysis setup. In a chi-squared ($\chi^{2}$) distribution with two degrees of freedom (d.o.f), the cumulative probability $P(TS>TS_{threshold})$ represents the likelihood of obtaining a test statistic above the threshold purely by chance. Using TS=12, the spurious detection probability is calculated as $\rm p\_value=1-\chi^{2}\times cdf(TS=12,\ d.o.f=2)\approx 0.0025$, where $cdf()$ is the cumulative distribution function. Given that 8200 seed positions were analysed, the expected number of spurious detections in the pre-selection phase is approximately $\rm 8200\times 0.0025=20.5$. This value represents an upper limit of contamination.
@@ 2.4 Additional Selection Criteria: Refining the 1CGH Sample
In the Fermi-LAT Pass 8 release (Atwood et al., 2013), each event is assigned a Point Spread Function (PSF) type, which indicates the quality of the reconstructed direction. Events are categorized into four quartiles: PSF0 (lowest quality, evtype=4), PSF1 (evtype=8), PSF2 (evtype=16), and PSF3 (highest quality, evtype=32). The PSF 68% containment radius varies by energy and event type. For example, at 30 GeV, the containment radius for PSF1 events is approximately 0.12, as described in the Pass 8 documentation [^4] Fermi-LAT Pass 8 PSF: https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm.
@@ 3 The First Cosmic Gamma-ray Horizon Catalogue (1CGH)
Figure 1: The distribution of detection significance ($\sigma$) for the 1CGH and 3FHL samples (respectively, blue line and red-dashed line).
In Fig. 1 we show the distribution of the detection significance for the entire 1CHG and 3FHL catalogues. This representation is meant as a qualitative overview on both samples, keeping in mind that the 1CGH is focus on seed-positions of blazars and blazar-candidates, while the 3FHL includes detections of all source types. The number of detections in 1CGH shows consistent growth along the significance-bins (i.e. following a trend similar to 3FHL), and includes an extra bin covering the 3<$\sigma$<4 range.
We should note that in the full 1CGH catalogue, 26 sources [^8] Comprising 3 unassociated sources, 4 BL Lacertae, 8 Flat Spectrum Radio Quasars (FSRQs), and 11 blazars of unknown type. reached the spectral index limit of $\Gamma=6.5$ constrained by our likelihood analysis setup (see Section 2.2). Moreover, 27 1CGH sources have counterparts in 3FHL where spectral curvature has been identified [^9] To help identify such cases, we include the Curvature_3FHL column to our final catalogue, which tracks whether spectral curvature was identified in 3FHL, and might affect broadband flux estimates.. For these 53 (26+27) cases, the derived fluxes should be interpreted as upper limits. While this affects only a small fraction of the sample, the vast majority of sources exhibit well-constrained spectral fits, reinforcing the robustness of the 1CGH catalogue for high-energy studies.
Figure 2: The photon index ($\Gamma$) versus the integral flux in the 10 GeV to 800 GeV energy band. The green points represent 1CGH sources with counterparts in both 3FHL and 4FGL-DR4 catalogues; the blue open diamonds represents 1CGH sources with counterparts in 4FGL-DR4 only; and the magenta circles represents new $\gamma$-ray identifications.
In Fig. 2 we present the photon index ($\Gamma$) versus the integrated flux derived from a power-law fit over the 10–800 GeV energy range. The plot shows that the 1CGH sample extends the 3FHL coverage toward fainter fluxes, effectively lowering the sensitivity limit for detections above 10 GeV. As in the 3FHL catalogue, the detection threshold in the 1CGH sample shows little dependence on photon index, with the new identifications clustering at the lowest fluxes. Moreover, the overlap between the 1CGH and 4FGL sources highlights a robust improvement in completeness above 10 GeV, which is complemented by the newly identified sources.
@@ 3.1 Extensive Redshift Review
Starting from the redshift information available in the reference catalogues (i.e., 5BZcat, 3HSP, 4LAC-DR3, and TeVcat), our literature review allowed us to assign or update redshift values for 967 1CGH sources –including 377, 437, and 153 sources with z-flags 1, 2, and 3, respectively. Of the total 2791 1CGH sources, 1062 (38.2%) have spectroscopically measured redshift values, 855 (30.6%) have uncertain/photometric values, 210 (7.5%) have lower-limit values, and 664 (23.7%) have no available redshift information.
@@ 3.2 The Cosmic Gamma-ray Horizon Plot
Here we draw attention to the number of 1CGH sources that lack robust redshift characterization; As mentioned in Section 3.1, a total of 1729 1CGH sources ($\sim$72%) are assigned redshift values that are either uncertain, lower limits, or are absent. All of those cases would benefit from dedicated observational campaigns to improve the overall redshift description, therefore representing a significant challenge for CGH studies and the very high-energy community as a whole.
@@ 3.3 Identification of Sources with Detectable EBL Absorption
In Figure 5, we illustrate the energy versus redshift relation and use the Saldana-Lopez EBL model Saldana-Lopez et al. (2021) as a reference framework to identify sources where gamma-ray absorption might be detectable. The Saldana-Lopez model offers an empirical determination of the evolving EBL spectrum and its uncertainties up to $\sim$6; it is based on galaxy counts and multifrequency data (from UV to far-IR) covering the Hubble Space Telescope’s ‘Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey’ (CANDELS), and was designed to minimise uncertainties especially at higher redshift (Grogin et al., 2011; Koekemoer et al., 2011).
However, our ultimate goal is to highlight a sample in which the $\tau_{(E,z)}$ values can be derived from the data, independently of any specific EBL model. The estimates of Saldana-Lopez et al. (2021) are similar to those by Finke et al. (2010) and both are representative of the most severe attenuation models among the ones considered.
We used an EBL optical depth of $\rm\tau_{(E,z)}=0.1$ as the lower limit to introduce an absorption flag (ABS-flag), indicating whether spectral absorption might be detectable at the highest energy bins. For cases where $\rm E^{bin}_{max}$ lies in the $\rm\tau_{(E,z)}>0.1$ regime, the largest energy bin observed with Fermi-LAT likely experiences a detectable absorption greater than 10% (i.e., >($\rm 1-e^{-0.1}$)). These sources can be used to investigate the opacity along redshift, and have ABS-flag set to ‘1’.
@@
The $\rm\tau_{(E,z)}$ values –calculated using the Saldana-Lopez EBL model at $\rm E^{bin}_{max}$ Saldana-Lopez et al. (2021)– are listed in the column ‘$\rm\tau_{Ebin}$’. Furthermore, the absorption fraction ($\rm 1-e^{-\tau_{(E,z)}}$) is recorded in the ‘$\rm Abs_{Ebin}$’ column, which can be used to create subsamples based on different absorption thresholds. In particular, for sources with multiple energy bins that experience detectable absorption, a single source can provide $\rm\tau_{(E,z)}$ estimates for multiple energy levels at a given redshift.
@@ 4 Summary and Conclusions
In this work, we presented the First Cosmic Gamma-ray Horizon (1CGH) catalogue, which lists all-sky Fermi-LAT detections above 10 GeV, compiled from 16 years of observations. The catalogue includes 2791 sources, with 62 representing new $\gamma$-ray detections, and to date, it represents a valuable sample to study the transparency of the Universe to VHE photons.
We significantly improved the redshift characterization of our catalogue by incorporating new spectroscopic redshift estimates whenever possible. By meticulously reviewing nearly 70 dedicated observational campaigns, we significantly reduced the redshift knowledge gap, introducing a z-flag system to categorise redshift reliability.
The 1CGH catalogue highlights sources experiencing varying degrees of $\gamma$-ray attenuation and will contribute to a better determination of the EBL’s density and its evolution over cosmic time. Using the Saldana-Lopez EBL model as a reference framework, we identified sources with $\gamma$-ray absorption above the $\rm\tau_{(E,z)}>0.1$ limit, creating a targeted sample where absorption is likely measurable.
In addition, we introduced a complementary data exploration method by calculating the mean energy of the four highest-energy photons $\rm E^{bin}_{max}$, rather than relying solely on the highest-energy event. This provides a more stable and statistically reliable measure of the characteristic ‘largest energy level’ observed for each source.
We identified sources that should be prioritised for follow-up optical observation campaigns. These include: (i) sources that lack redshift information but have a clear optical or radio association, prioritised based on the highest $\rm E^{bin}_{max}$ values; and (ii) sources with lower limit or uncertain redshift estimates that could benefit from further spectroscopic exploration, prioritised according to their $\tau$ values measured at $\rm E^{bin}_{max}$. These targets will not only help refine the redshift description of the catalogue but also improve the accuracy of EBL measurements at higher redshift, thus reducing uncertainties in SFR density determination during the EoR, at $\sim$6-7.
Following our redshift review (Section 3.1), approximately 72% of the 1CGH sources still lack robust redshift determination. This significant gap poses challenges for the very high-energy community as a whole, and calls attention to the need for extensive follow-up optical observation campaigns targeting $\gamma$-ray blazars.
@@ Appendix A Column Descriptions of the 1CGH Catalogue
The 1CGH catalogue contains detailed metadata for 2791 sources detected above 10 GeV based on 16 years of Fermi-LAT observations. Table 2 provides a description of each column and details its content.
@@
zflag_3HSP Redshift flag from the 3HSP catalogue: (1) spectroscopic, (2) uncertain, (3) lower limit redshift, (4) photometric -with featureless optical spectrum, (5) photometric -without optical spectrum.